If a bank uses $80 of reserves to make a new loan when the reserve ratio is 25%

A) Money supply initially decreases by $80

B) Money supply initially increases by $20

C) Money supply will eventually increase by more than $20 but less than $80

D) Level of wealth in economy will not change
Answer #1

Ans: Level of wealth in economy will not change.

Explanation:

Money multiplier = 1 / RR = 1 / 0.25 = 4

Increase in money supply = $80 * 4 = $320

But, increases in money supply does not lead to increase in wealth.

Thus, option [D] is correct answer.